# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 14.10 Type ${\bf A}_{2}$
Let $\omega_{1}$ and $\omega_{2}$ be the two fundamental weights for ${\bf SL}_{3}$ and $\alpha_{1},\alpha_{2}$ the simple roots. As before, set $X:={\bf SL}_{3}/{B}$ and $Y_{i}:={\bf SL}_{3}/{P}_{\alpha_{i}},i=1,2$.
@@
Proof.
Proposition 14.18 and Theorem 14.1 give ${\sf A}_{0}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=1,2,3$. Proposition 14.23 gives ${\sf A}_{-1}\subset{{}^{\perp}{\sf A}_{-2}}$. Finally, Proposition 14.10 gives ${\sf A}_{-i}\subset{{}^{\perp}{\sf A}_{-3}}$ for $i=0,1,2$. This proves that the sequence $\langle{\sf A}_{-3},\dots,{\sf A}_{0}\rangle$ is semiorthogonal.
By Corollary 14.4, the terms outside the block ${\sf A}_{-2}$ are exceptional: these are either line bundles or the bundles in the block ${\sf A}_{-1}$ which are obtained by the left mutation of the block $\mathcal{L}(\omega_{1}),\mathcal{L}(\omega_{2})$ through $\mathcal{L}(0)$; since the latter block is exceptional, then so is the mutated block ${\sf A}_{-1}$. Proposition 14.25 states that the bundles ${\sf D}_{\alpha_{i}}\mathcal{L}(\omega_{i})$ are exceptional for $i=1,2$ and so are the bundles ${\sf D}_{\alpha_{i}}(\Psi_{1}^{\omega_{i}})$, their left mutations through the exceptional bundle $\mathcal{L}(0)$. Proposition 14.22 allows to identify the bundles ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ and ${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})$; it is sufficient to compute one of these, say ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ that is the kernel of the map $\Psi_{1}^{\omega_{1}}\rightarrow\mathcal{L}(\omega_{1}-\alpha_{1})$. The weights of $\Psi_{1}^{\omega_{1}}$ are $\omega_{1}-\alpha_{1}$ and $-\omega_{2}$, thus the kernel ${\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})$ is isomorphic to $\mathcal{L}(-\omega_{2})$. Similarly, ${\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})=\mathcal{L}(-\omega_{1})$. This implies $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}),{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))=\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}),{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))=0$. Thus, ${\sf A}_{-2}$ is an exceptional block. Finally, Proposition 14.27 implies that the minimal full triangulated subcategory of $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$ that contains all ${\sf A}_{-i},i=0,\dots,3$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(X)$. ∎
@@
Proof.
According to Proposition (7.5), one needs to check the following orthogonalities:
  1. [1)]
  2. 1.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-3],\mathcal{L}(-\rho))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(\rho)[-3],{{\sf A}_{i}})=0$ for $i\neq-3$
  3. 2.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{1}}\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  4. 3.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}}))=0$;
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\Psi_{1}^{\omega_{2}}\otimes\mathcal{L}(\rho)[-2],{{\sf A}_{i}})=0$ for $i\neq-2$
  5. 4.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{1}}))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{2}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  6. 5.
    $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{2}}))={\sf k}$; $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],\Psi_{1}^{\omega_{1}}))=0$;  $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\otimes\mathcal{L}(\rho)[-1],{{\sf A}_{i}})=0$ for $i\neq-1$
  7. 6.
    $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),\mathcal{L}(0))={\sf k}$;  $\mathop{\mathrm{RHom}}\nolimits_{X}(\mathcal{L}(0),{{\sf A}_{i}})=0$ for $i\neq 0$.
[…]
@@ 14.11 The quiver for ${\bf A}_{2}$
The picture below depicts the path quiver algebra of the above full exceptional collection. There are only zero-degree $\mathop{\mathrm{Hom}}\nolimits$-spaces in the non-trivial direction in the collection (76). The underlying graph of the quiver is the Bruhat graph for $SL_{3}/{B}$; the quiver path algebra is generated in degree one and the relations are quadratic that one reads off the $\mathop{\mathrm{Hom}}\nolimits$–spaces next to the arrows of the quiver. The labels assume $p$ is not too small. For small $p$ the characters are still correct and the $\mathop{\mathrm{Hom}}\nolimits$–spaces have a good filtration.
${{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\Psi_{1}^{\omega_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$${\mathcal{L}(-\rho)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\quad}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{1}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{1}+\rho})}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{2}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{2}+\rho})}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\quad\mathcal{L}(0)}$${{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$$\scriptstyle{\nabla_{\omega_{1}}}$${\Psi_{1}^{\omega_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$ (78)
The Weyl group elements that we associate to the nodes are as follows: The leftmost node belongs with the longest element $w_{0}$, the rightmost with the identity element. In the top row we have from left to right $s_{2}s_{1}$, $s_{2}$, and in the bottom row from left to right $s_{1}s_{2}$, $s_{1}$.
@@ 14.12 Type ${\bf B}_{2}$
We assume here that the prime $2$ is invertible in ${\sf k}$. Let $G$ be a simple group of type ${B}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${B}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${P}_{\alpha}$ (resp., ${P}_{\beta}$) be the minimal parabolic generated by ${B}$ and $U_{\alpha}$ (resp., ${B}$ and $U_{\beta}$). The Grassmannian ${G}/{P}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=5$. The Grassmannian ${G}/{P}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=4$ (it is in fact isomorphic to $\mathbb{P}^{3}=\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast}))$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (56).
@@
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/{B})$:
@@
The $G$-equivariant vector bundles in the above collection can be explicitly identified:
@@
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{G/B}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(G/B,\mathcal{L}(-\beta))=k$. The bundle $\Psi_{2}^{\omega_{\alpha}}$ is the left mutation of line bundle $\mathcal{L}(2\omega_{\alpha})$ through the admissible subcategory $\langle\mathcal{L}(0),\mathcal{L}(\omega_{\alpha})\rangle\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/B)$ and fits into a short exact sequence
@@
is a full exceptional collection on $G/P_{\beta}$. To ensure that ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}=\Psi_{2}^{\omega_{\alpha}}$ it is sufficient to check that $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(0),{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$, $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$, and the ranks of bundles in Remark 14.34, (2) are the same. That the ranks are the same is immediate from Proposition 14.21. As for the orthogonality, the first equality holds by Proposition 14.18, the second one by Proposition 14.23, and the third one is a consequence of $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$. The latter equality is obtained by applying $\mathop{\mathrm{Hom}}\nolimits_{X}(-,\mathcal{L}(-\omega_{\alpha}))$ to the sequence (64) and then using Proposition 14.18, the exact sequence (47), and Theorem 14.1: all weights that are needed to compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\mathcal{L}(\omega_{\beta})),\mathcal{L}(-\omega_{\alpha}))$ are singular.
Proof.
Theorem 14.1 and Proposition 14.18 give ${\sf A}_{0}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=1,\dots,4$. Proposition 14.23 and Proposition 14.10 give ${\sf A}_{-1}\subset{{}^{\perp}{\sf A}_{-i}}$ for $i=2,\dots,4$. By Proposition 14.10 ${\sf A}_{-i}\subset{{}^{\perp}{\sf A}_{-4}}$ for $i=2,3$. Showing ${\sf A}_{-2}\subset{{}^{\perp}{\sf A}_{-3}}$ amounts to the vanishing
$\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0,\ \mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$ (83)
$\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=0,\ \mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=0$ (84)
By Proposition 14.8, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(-\alpha)[1])$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},{\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\beta}})\otimes\mathcal{L}(-\beta)[1])$. It follows from Remark 14.34 that ${\sf D}_{\alpha\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})=0$ and ${\sf D}_{\beta\alpha\beta\alpha}(\Psi_{1}^{\omega_{\beta}})=0$ as $\langle-\omega_{\alpha},\alpha^{\vee}\rangle=\langle-\omega_{\beta},\beta^{\vee}\rangle=-1$. Hence, the groups from (83) are trivial.
By Proposition 14.7, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))$. Let us show that both groups are trivial; it proves at the same time that ${\sf A}_{-2}$ is a block. By Proposition 14.8, there are isomorphisms $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\otimes\mathcal{L}(-\beta)[1])$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}))=\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))\otimes\mathcal{L}(-\alpha)[1])$. By Remark 14.34, ${\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})$ and ${\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}}))$ are identified with line bundles $\mathcal{L}(-\omega_{\alpha})$ and $\mathcal{L}(-\omega_{\beta})$, respectively. Thus, we need to compute the groups $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}-\beta)[1])$ and $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},\mathcal{L}(-\omega_{\beta}-\alpha)[1])$. Using the sequence (56), this amounts to computing cohomology of line bundles; one easily checks that all the weights occurring are singular and thus all the groups vanish by Theorem 14.1. This proves the vanishing of groups in (84) and ${\sf A}_{-2}$ being a block.
Block structure in blocks ${\sf A}_{-1},{\sf A}_{-3}$ follows easily again from Theorem 14.1. Note that by [8], the collection of bundles $\langle\mathcal{L}(0),\Psi_{1}^{\omega_{\beta}},{\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}}),{\sf D}_{\alpha\beta}(\Psi_{1}^{\omega_{\beta}})\rangle$ generates the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P_{\alpha})$ and the collection of bundles $\langle\mathcal{L}(0),\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}}),{\sf D}_{\beta\alpha}(\Psi_{1}^{\omega_{\alpha}})\rangle$ generates the category $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(G/P_{\beta})$. By Proposition 14.27, the collection of subcategories ${\sf A}_{-i},i=0,\dots 4$ is a semiorthogonal decomposition. Finally, all the line bundles in (79) are exceptional by Corollary 14.4, while ${\sf D}_{\beta}(\Psi_{1}^{\omega_{\beta}})$ is exceptional by Remark 14.34, (2) and ${\sf D}_{\alpha}(\Psi_{1}^{\omega_{\alpha}})$ is exceptional by Corollary 14.26. ∎

### v3

@@ 14.10 Type ${\bf A}_{2}$
Let $\omega_{1}$ and $\omega_{2}$ be the two fundamental weights for ${\bf SL}_{3}$ and $\alpha_{1},\alpha_{2}$ the simple roots. As before, set $X:={\bf SL}_{3}/{\mathbf{B}}$ and $Y_{i}:={\bf SL}_{3}/{\mathbf{P}}_{\alpha_{i}},i=1,2$.
@@ 14.11 The quiver for ${\bf A}_{2}$
The picture below depicts the path quiver algebra of the above full exceptional collection. There are only zero-degree $\mathop{\mathrm{Hom}}\nolimits$-spaces in the non-trivial direction in the collection (84). The underlying graph of the quiver is the Bruhat graph for $SL_{3}/{\mathbf{B}}$; the quiver path algebra is generated in degree one and the relations are quadratic that one reads off the $\mathop{\mathrm{Hom}}\nolimits$–spaces next to the arrows of the quiver. The labels assume $p$ is not too small. For small $p$ the characters are still correct and the $\mathop{\mathrm{Hom}}\nolimits$–spaces have a good filtration.
${{\sf D}_{\alpha_{1}}(\Psi_{1}^{\omega_{1}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\Psi_{1}^{\omega_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$${\mathcal{L}(-\rho)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\quad}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{1}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{1}+\rho})}$$\scriptstyle{{\sf Ker}(\nabla_{\omega_{2}}\otimes\nabla_{\rho}\rightarrow\nabla_{\omega_{2}+\rho})}$$\scriptstyle{\nabla_{\omega_{1}}}$$\scriptstyle{\nabla_{\omega_{2}}}$${\quad\mathcal{L}(0)}$${{\sf D}_{\alpha_{2}}(\Psi_{1}^{\omega_{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{2}}}$$\scriptstyle{\nabla_{\omega_{1}}}$${\Psi_{1}^{\omega_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nabla_{\omega_{1}}}$
The Weyl group elements that we associate to the nodes are as follows: The leftmost node belongs with the longest element $w_{0}$, the rightmost with the identity element. In the top row we have from left to right $s_{2}s_{1}$, $s_{1}$, and in the bottom row from left to right $s_{1}s_{2}$, $s_{2}$.
@@ 14.12 Type ${\bf B}_{2}$
We assume here that the prime $2$ is invertible in ${\sf k}$. Let $\mathbf{G}$ be a simple group of type ${\mathbf{B}}_{2}$. Let $\alpha$ and $\beta$ be the simple roots for ${\mathbf{B}}_{2}$ with $\beta$ being the long root. Let $\omega_{\alpha}$ and $\omega_{\beta}$ be the two fundamental weights. Let ${\mathbf{P}}_{\alpha}$ (resp., ${\mathbf{P}}_{\beta}$) be the minimal parabolic generated by ${\mathbf{B}}$ and $\mathbf{U}_{\alpha}$ (resp., ${\mathbf{B}}$ and $\mathbf{U}_{\beta}$). The Grassmannian ${\mathbf{G}}/{\mathbf{P}}_{\alpha}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\beta}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\beta})$ and ${\rm dim}\nabla_{\omega_{\beta}}=5$. The Grassmannian ${\mathbf{G}}/{\mathbf{P}}_{\beta}$ is embedded into $\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast})$ with the help of line bundle $\mathcal{L}(\omega_{\alpha})$ and ${\rm dim}\nabla_{\omega_{\alpha}}=4$ (it is in fact isomorphic to $\mathbb{P}^{3}=\mathbb{P}(\nabla_{\omega_{\alpha}}^{\ast}))$. Associated to the fundamental weights $\omega_{\alpha}$ and $\omega_{\beta}$ are the vector bundles $\Psi_{1}^{\omega_{\alpha}}$ and $\Psi_{1}^{\omega_{\beta}}$ defined as in the sequence (64).
@@
The following collection is a full exceptional collection in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/{\mathbf{B}})$:
@@
The $\mathbf{G}$-equivariant vector bundles in the above collection can be explicitly identified:
@@
the corresponding extension is given by the isomorphism ${\mathrm{Ext}}^{1}_{\mathbf{G}/\mathbf{B}}(\mathcal{L}(\omega_{\alpha}-\omega_{\beta}),\mathcal{L}(-\omega_{\alpha}))={\rm H}^{1}(\mathbf{G}/\mathbf{B},\mathcal{L}(-\beta))=k$. The bundle $\Psi_{2}^{\omega_{\alpha}}$ is the left mutation of line bundle $\mathcal{L}(2\omega_{\alpha})$ through the admissible subcategory $\langle\mathcal{L}(0),\mathcal{L}(\omega_{\alpha})\rangle\subset\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}(\mathbf{G}/\mathbf{B})$ and fits into a short exact sequence
@@
is a full exceptional collection on $\mathbf{G}/{\mathbf{P}}_{\beta}$. To ensure that ${\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}}=\Psi_{2}^{\omega_{\alpha}}$ it is sufficient to check that $\mathop{\mathrm{RHom}}\nolimits(\mathcal{L}(0),{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$, $\mathop{\mathrm{RHom}}\nolimits(\Psi_{1}^{\omega_{\alpha}},{\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}})=0$ and $\mathop{\mathrm{RHom}}\nolimits({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$, and the ranks of bundles in Remark 14.35, (2) are the same. That the ranks are the same is immediate from Proposition 14.22. As for the orthogonality, the first equality holds by Proposition 14.19, the second one by Proposition 14.24, and the third one is a consequence of $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}\Psi_{1}^{\omega_{\beta}},\mathcal{L}(-\omega_{\alpha}))=0$. The latter equality is obtained by applying $\mathop{\mathrm{Hom}}\nolimits_{X}(-,\mathcal{L}(-\omega_{\alpha}))$ to the sequence (72) and then using Proposition 14.19, the exact sequence (57), and Theorem 14.1: all weights that are needed to compute $\mathop{\mathrm{RHom}}\nolimits_{X}({\sf D}_{\beta}(\mathcal{L}(\omega_{\beta})),\mathcal{L}(-\omega_{\alpha}))$ are singular.